OpenStudy (voltage):

Find the value of the following expression: (38 ⋅ 2−5 ⋅ 90)−2 ⋅ 2 to the power of negative 2 over 3 to the power of 3, whole to the power of 4 ⋅ 328 (5 Write your answer in simplified form. Show all of your steps

OpenStudy (hoot_hoot):

Is this your question: \[[(38 \times 2 - 5 \times 90) - 2 \times \frac{2^-2 }{ 3^3 } ]^4 \times 328 (5)\]

OpenStudy (hoot_hoot):

\[[(76 - 5 \times 90) - 2 \times \frac{2^2}{3^3}]^4 \times 1640 = [(76 - 450) - \frac{2^3}{3^3}]^4 \times 1640 = [-374 - \frac{8}{27}]^4 = [\frac{ -10106 }{ 27 }]^4 \times 1640 = 19 627 370 427 \times 1640 = 3.21889E+13\]

OpenStudy (hoot_hoot):

If the given is correct...